OpenStudy (migitmack):

Can some one explain 2x^2-3x=0

Parth (parthkohli):

Hint: Use factoring.

Parth (parthkohli):

Can you factor \(2x^2 - 3x\)?

OpenStudy (migitmack):

Ughm Yes

Parth (parthkohli):

So please do.

OpenStudy (migitmack):

Does it become 2x and 3x

OpenStudy (migitmack):

I am not to sure

OpenStudy (migitmack):

@ParthKohli

Parth (parthkohli):

The \(x\) gets out of the bracket.

Parth (parthkohli):

\[x(2x - 3) = 0 \] right?

OpenStudy (migitmack):

yes

Parth (parthkohli):

Ok. Now use the zero-product rule.

Parth (parthkohli):

\[x(2x - 3) = 0 \Longrightarrow x =0, 2x - 3 =0 \] Can you do it now?

OpenStudy (migitmack):

mhm